OpenStudy (anonymous):

If someone has a networth of 7.92×10^10 .Suppose that this individual now decides to give one-eighth of a percent of his new net worth to charity. How many dollars are given to charity?

OpenStudy (anonymous):

1/8 in decimal form is (1/8)/100 = .00125 .00125 * 7.92×10^10 = 9.9*10^7

OpenStudy (anonymous):

what does 1/8th of a percent mean

OpenStudy (anonymous):

1% is .01. divide that by 8

OpenStudy (anonymous):

oh okay ty